# Parallel Scan on Ascend AI Accelerators

### v1

@@ 4.2. Multiple (Batched) Scans
Let’s discuss, as an example, a particular case where ScanU is used as a building block. The batched algorithm uses the same principles as the Algorithm 1 but also considers the 2-to-1 ratio between the vector and cube cores in the split Ascend architecture (910B). Figure 4 depicts the main ideas behind our batched scan algorithm in this case. The algorithm starts by computing the local scans of size $s$ of $\bm{x}$ of all input arrays. Each cube core computes the local scans of a tile of size $\ell$ in two batches at the same time; once the tiles of the first two batches are ready, two distinct vector cores will complete the scans independently over each batch by propagating the partial sums within the tiles. This process is pipelined through AscendC to efficiently use all available hardware (vector) resources. The second batched scan algorithm extends ScanUL1 (Algorithm 2) so that each AI core computes a scan on a separate array in the batch.
@@
Figure 5. Execution time ratio between ScanUL1 and ScanU batched scan algorithms for various array length ($x$-axis) and batch sizes ($y$-axis). Baseline is ScanU.
[…]
[Execution time ratio between ScanUL1 and ScanU.]Execution time ratio between ScanUL1 and ScanU batched scan algorithms for various array length.
@@
Figure 5 compares the two batched scan algorithms presented for varying input batch size and array length. The first batched scan algorithm, based on ScanU (Algorithm 1), is used as our reference/baseline. Both algorithms have the same tiling strategy based on the input shapes to ensure a fair comparison. The figure demonstrates that the algorithms perform well in different cases and, more importantly, complement each other. In particular, ScanU is superior when the batch size is greater than $18$, and the input length is smaller than $4K$. On the other hand, ScanUL1 is superior when the batch size is smaller than $18$ and the input length larger than $4K$.
4.3. Multi-core Scan
This section presents a multi-core scan algorithm (MCScan, Algorithm 3) to compute the prefix sum of large input arrays. For the sake of a clearer explanation, the presented algorithm uses a vector-to-cube ratio of 1-to-1. However, our implementation takes advantage of the 2-to-1 ratio of 910B, but we consider it an implementation detail.
MCScan is similar to the Scan-Scan-Add (SSA) paradigm of computing scans using a hierarchical partition of the input into blocks (top-level of the hierarchy) and tiles within each block as discussed in Section 2.1. MCScan consists of two phases separated by a global synchronization barrier across all cores (blocks). Figure 6 depicts these two phases of the algorithm. In contrast to previous methods, MCScan performs partial re-computation of the block-level reductions during its first phase, as we will explain shortly.
In the first phase, the cube and vector units work in parallel to partially compute the first scan part of SSA simultaneously. The cube units compute the local prefix sums of all consecutive tiles of size $s$ and write them back to global memory. In parallel with the cube units, the vector units compute the reduction over the tiles and then hierarchically reduce the tile reductions on a block-level granularity. The result of these reductions is written in an array $\bm{r}$ where the $i$-th entry equals the sum of all the values of the $i$-th block. By definition, $\bm{r}$ has length equal to the number of blocks, $B$.
1:procedure MCScan($\bm{x}$, $s$, $B$) $\triangleright$ $B$: number of blocks
2:     Let $\bm{y}$ be the output array
3:     Let $\bm{r}$ be an array of length $B$ in $\mathtt{GM}$
4:     parfor $i$-th block of $\bm{x}$: $\bm{x}[i]$ $\triangleright$ Phase I
5:         Load $\bm{U}_{s}$ in $\mathtt{L0B}$triangleright Cube Units
6:         for each $\ell$-tile of $\bm{x}[i]$: $\bm{x}_{\ell}$ do
7:              Load $\bm{x}_{\ell}$ from $\mathtt{GM}$ to $\mathtt{L0A}$$\triangleright$ Cube Units
8:              $\bm{C}\leftarrow\bm{A}_{s}\ @\ \bm{U}_{s}$triangleright Cube Units
9:              Copy $\bm{C}$ in $\bm{y}[i]$ in $\mathtt{GM}$triangleright Cube Units
10:         end for
11:         Load $\bm{x}[i]$ to $\mathtt{UB}$triangleright Vector Units
12:         $r_{i}\leftarrow$ ReduceSum($\bm{x}[i]$) $\triangleright$ Vector Units
13:         Write $r_{i}$ on $i$-th entry of $\bm{r}$ in $\mathtt{GM}$triangleright Vector Units
14:     end parfor
15:     SyncAll: Synchronize all cube/vector cores
16:     parfor $i$-th block of $\bm{y}$: $\bm{y}[i]$ do $\triangleright$ Phase II
17:         Load $\bm{r}$ from $\mathtt{GM}$ to $\mathtt{UB}$
[…]
19:         for each $\ell$-tile of $\bm{y}[i]$: $\bm{y}_{\ell}$ do
20:              for each $s$-tile of $\bm{y}_{\ell}$: $\bm{y}_{s}$ do
21:                  $\bm{y}_{s}\leftarrow\bm{y}_{s}+\textit{partial}$triangleright In-place add
[…]
23:              end for
24:              Copy $\bm{y}_{\ell}$ from $\mathtt{UB}$ to $\mathtt{GM}$
25:         end for
26:     end parfor
27:     Return $\bm{y}$
28:end procedure
Algorithm 3 Multi-core Scan (MCScan)
Figure 6. Multi-core scan (Algorithm 3) consists of two phases synchronized by a global barrier. The first phase computes tile-level (size $s$) local scans on cube units and block-level reductions in vector units. The second phase reads local scans and reductions from global memory to the vector cores where the scan computation is completed.
[…]
[Multi-core scan (Algorithm 3) consists of two phases synchronized by a global barrier.]Multi-core scan (Algorithm 3) consists of two phases synchronized by a global barrier.
Exclusive scan and int8 support
@@ 4.3. Multi-core Scan
Here, we discuss a few extensions of the multi-core scan algorithm that we have implemented. In particular, we added support for exclusive scans and integer inputs. Typically, AI accelerators support low-precision arithmetic since inference of deep learning models is robust to extreme levels of quantization (Lin et al., 2024). In particular, Ascend supports input matrices of 8-bit integer data type with output/accumulation in 32-bit integer data type. Since the scan is a memory-bound operator, there is an opportunity to improve performance in terms of elements per second processed; see Figure 9. We have implemented a specialization of Algorithm 3 for integers with 8 bits.
@@
We implemented exclusive scan by writing the output of inclusive scan to global memory shifted by one element, discarding the last value and writing zero to the first position by a single block.
5. Operators based on Scan
In this section, we revisit several computational parallel primitives based on parallel scans (Blelloch, 1989; Sengupta et al., 2007). Scan-based primitive include weighted sampling, split and compress/compact (equivalent to the masked_select Pytorch operator). It is well-known that radix sort can be implemented on top of split (Blelloch, 1990, Section 1.3). Also top-k can be implemented on top of split using a partial quick-sort/select approach (Blelloch, 1996).
Interestingly enough, current AI workloads like Large Language Model (LLM) inference make implicitly heavy use of scan-based computational primitives, including top-$k$ and top-$p$ (nucleus) sampling (Holtzman et al., 2020), see also (Gu and Dao, 2024). The top-$p$ sampling implementation of the popular open-source model Llama3 (Touvron et al., 2023) contains a batched sorting and prefix sum operation as the first two PyTorch operations, see (Meta AI, 2024).
Figure 7. A diagram of well-known parallel scan applications considered here along with their dependencies.
[…]
[A diagram of parallel scan applications considered here along with their dependencies.]A diagram of parallel scan applications considered here along with their dependencies.
In the rest of the section, we describe the split, compress, radix sort, top-$k$, top-$p$ sampling and weighted sampling operators in more detail. Figure 7 depicts the dependencies between the scan-based primitives. Top-$p$ sampling extensively uses multiple scan operations, as discussed in Section 5.
Split
Compress
Radix sort
Radix sort is a well-known application of the split operator (Blelloch, 1990; Blelloch et al., 1991). A radix sort algorithm loops over the bits of the input elements, starting at the least significant bit and executes a split where the mask is obtained by reading the corresponding bit (radix) on each iteration. We implement a Least-Significant Bit (LSB) radix sort in AscendC using the split operator based on the MCScan algorithm. We implemented an additional vector-only kernel, RadixSingle, that extracts the radices of the inputs before the execution of the split. RadixSingle makes use of the AscendC vector instructions ShiftRight and Not to create the input mask for split. Additional pre-processing and post-processing phases are needed to support floats; see Exercises 8 and 9 in (Knuth, 1998, Section 5.2.5). The pre-processing phase encodes all the input elements by inverting the Most Significant Bit (MSB) of positive numbers and all the bits of the negative numbers. Applying an unsigned integer radix sort on the encoded elements will correctly order them. The post-processing phase is needed to decode the elements back to the original value. We have implemented the pre- and post-processing steps using AscendC bit-wise vector instructions, and thus, we support sorting of fp16 data types.
The paper on evaluating radix sort using the Connection Machine (CM-2) came to our attention in the later stages of our radix sort development (Blelloch et al., 1991). In this work, the authors share several important implementation details that were not known to us a priori. For example, the fact that radix sort works with floats is interesting and is quite useful when working with low bit-width floats.
Top-$k$
Top-$k$ selection is an essential operation in various settings, including similarity search queries (Johnson et al., 2021) and Large Language Models (LLMs) inference where the output tokens are typically sampled from the $k$ tokens having the highest probability for the given context (Kool et al., 2019). The interested reader is referred to a recent survey on parallel top-$k$ (Zhang et al., 2023). A recent work on top-$k$ is Radik, a Radix-based GPU implementation that scales well for large values of $k$ (Li et al., 2024a, b).
Top-$p$ or nucleus Sampling
Top-$p$ sampling in Large Language Model inference is an additional operation that applies sort and scan on the token probability vector (Holtzman et al., 2020). These operations are usually batched with a constant batch size. Interestingly, if the sorting step is implemented using radix sort, the top-$p$ sampling operator becomes a scan-intensive operator! Indeed, top-$p$ executes 17 scans for each batch: 16 scan operations for radix sort (one scan per bit, fp16) and an additional scan required by the algorithm.
Weighted Sampling
@@ 5. Operators based on Scan
We implement a parallel weighted sampling kernel using the well-known inverse transform sampling approach and a scan to compute the cumulative distribution. Given an array $\bm{w}$ of $n$ positive weights, the goal is to draw a sample with proportional probability to the weights. The output is an index $i$ of $\bm{w}$ with probability proportional to $w_{i}$. First, we scan $\bm{w}$, and then, given a uniform sample $\theta\in{[0,1]}$, we invoke the SplitInd kernel with input $\text{scan}(\bm{w})$ and the element-wise predicate ?$>\theta*\sum_{i}{w_{i}}$. The last entry of the output indices array of SplitInd contains the weighted sample. For more advanced parallel weighted sampling techniques, see (Hübschle-Schneider and Sanders, 2022).
@@
The performance improvement of our proposed weighted sampling kernel is not significant compared to the baseline for a single sample. However, our implementation does provide a functional improvement compared to the baseline operator. Indeed, the baseline Ascend weighted sampling operator torch.multinomial supports discrete distributions with support size up to $2^{24}$ elements, whereas our approach can support distributions with arbitrary support size. We leave as future work any further possible improvements on parallel weighted sampling. In particular, for the multiple sample generation scenario, the parallel alias table construction of (Hübschle-Schneider and Sanders, 2022) seems to be a promising direction.
6. Experimental Evaluation
We evaluate the multi-core scan, compress, radix sort, batch scan, and top-$p$ sampling. We have implemented all proposed algorithms in C++17 using the AscendC programming framework discussed in Section 3.2. We used the Ascend CANN toolkit 8.0.RC3.alpha002 with Ascend firmware and drivers versions 1.0 and 23.0.0, respectively. All evaluations are performed on Huawei’s Ascend 910B4 accelerator. In particular, 910B4 contains 20 Cube Units and 40 Vector Units (the vector-to-cube units ratio is 2-to-1). The host CPU is an AMD EPYC processor running on Ubuntu 22.04.
All timing measurements are collected using the PyTorch profiler functionality in Python. We used Ascend’s PyTorch adapter [^2] https://gitee.com/ascend/pytorch with version v2.1.0 to report all our PyTorch-related measurements. To wrap our custom AscendC operators in PyTorch, we used the open-sourced operator plugin framework at https://gitee.com/ascend/op-plugin. op-plugin allows its users to easily define a custom PyTorch operator. Before wrapping our operators using the op-plugin, we used the msopgen [^3] https://www.hiascend.com/document/detail/en/canncommercial/700/operatordev/tbeaicpudevg/atlasopdev_10_0024.html tool to wrap the AscendC operators.
Evaluation against other accelerators/architectures
6.1. Multi-core Scan (MCScan)
Figure 8. Bandwidth of MCScan (Algorithm 3) for $s=32,64,128$. In addition, the copy operator (torch.clone) is depicted for comparison. Memory bandwidth (910B4) is $800$ GB/s.
[…]
[Bandwidth of multi-core scan]Bandwidth of multi-core scan
Figure 8 depicts the performance of MCScan (Algorithm 3) on Ascend 910B4 versus the state-of-the-art (torch.cumsum). We integrated the multi-core algorithm in PyTorch to ensure a fair comparison and expose it as a custom PyTorch operator. The custom PyTorch operator statically pre-allocates an upper triangular all-ones matrix $\bm{U}_{s}$ for all $s=32,64,128$. The baseline operator doesn’t use the cube unit, while MCScan takes advantage of all the computing units reaching up to 37.5% of theoretical memory bandwidth (peak bandwidth is 800GB/s for 910B4). To have a more solid evaluation of our implementation, we compare it to a copy kernel that performs a memory copy; we used the torch.clone(). This shows us that for sizes smaller than the L2 cache, we almost approach the theoretical limit given by the memory bandwidth. A clear trend is that the larger the matrix multiplication dimension $s$ is, the better the performance of the multi-core scan. $s=128$ maximises the utilization of the level-0 scratchpad memories $\mathtt{L0A}$ and $\mathtt{L0B}$ of the cube unit. We have measured the speed-up between MCScan and ScanUon 910B4 (20 AI cores), and it saturates at 15.2×15.2$\times$15.2 × for large input sizes. We foresee that increasing the matrix multiplication tile size could lead to further performance improvements, but we leave this investigation as future work.
@@ 6.1. Multi-core Scan (MCScan)
Next, we investigate the additional performance benefits of taking advantage of the lower-precision input data (int8) capability of the cube unit. Figure 9 depicts the performance of the multi-core scan algorithm in terms of giga elements per second for input data types float16 and $8$-bit integers (int8). As it is depicted, there is a performance improvement of the order of 10% for the case of integer inputs. Such an improvement is crucial since the split and compress operators take as input boolean mask arrays that are stored [^4] This is due to the PyTorch framework, see for example https://github.com/pytorch/pytorch/issues/41571. in int8 format.
@@
Figure 9. Giga elements per second comparison of MCScan for float16 (fp16) and 8-bit integers (int8) input data types.
[…]
[Giga elements per second comparison of MCScan for float16 (fp16) and 8-bit integers (int8) input data types.]Giga elements per second comparison of MCScan for float16 (fp16) and 8-bit integers (int8) input data types.
6.2. Compress
Figure 10. Bandwidth of the compress operator based on MCScan ($s=32,64,128$) and the baseline torch.masked_select operator.
[…]
[Bandwidth of the compress operator.]Bandwidth of the compress operator.
Figure 10 depicts the performance comparison between compress versus the baseline PyTorch masked_select. We set the mask so that each mask entry is independently set to true or false uniformly at random. The figure indicates that the baseline masked_select operator is not optimized on Ascend, and a code investigation reveals that the baseline does not use the vector or cube units. On the other hand, our Compress kernel reaches up to 160GB/s (20% of peak memory bandwidth).
6.3. Radix sort
We modified the radix sort operator to additionally return indices that correspond to the input index of each output element. This modification ensures a fair comparison with the sort operator provided by the PyTorch Ascend adapter. The modification is based on the SplitInd operator by keeping track of the output indices on each split application. Our radix sort implementation is stable and supports unsigned (or signed) integers and floats (fp16).
Figure 11. Comparison between radix sort and torch.sort() for floating-point numbers in half-precision ($16$ bits).
[…]
[Comparison between radix sort and torch.sort() for floating-point numbers in half-precision ($16$ bits).]Comparison between radix sort and torch.sort() for floating-point numbers in half-precision ($16$ bits).
Figure 11 depicts the performance of a parallel fp16 radix sort implementation using MCScan with input data type int8 (Algorithm 3) to perform the parallel split step. For input lengths greater than 525K, our “textbook” implementation of radix sort delivers a speedup between 1.3×1.3$\times$1.3 × up to $\times3$ compared to the torch.sort() baseline.
We expect additional performance improvements for low-precision data types (low bit-width) since the number of radix sort iterations equals the input bit-width. Indeed, the trend in AI accelerators is to introduce low-precision formats like $8$-bit floats or $4$-bit integers (int4) (NVIDIA Corporation, 2020). Therefore, an additional performance improvement (2×2$\times$2 ×) for sorting in low-precision $8$-bit scenarios is expected without further development effort.
6.4. Batched Scan
Figure 12 depicts the performance of the batched scan kernels on Ascend for varying input batch sizes and input lengths equal to $65K$. We depict the memory bandwidth achieved for tiling parameters $s=16,32,64$ and $128$. Our proposed batch scan operators for $s=64$ and $128$ reach up to $400$ GB/s. Interestingly enough, for smaller values of $s=16,32$, the performance of the proposed batch scan kernels is poor. In addition, the performance of our proposed batch scan kernel for $s=16$ and the baseline is similar.
Figure 12. Bandwidth of batched scan based on Algorithm 1 for increasing batch sizes and $s=16,32,64,128$. Input length is $65K$. Memory bandwidth (910B4) is $800$ GB/s.
[…]
[Bandwidth of batched scan.]Bandwidth of batched scan.
6.5. Top-$p$ Sampling
Figure 13. Execution time (in milliseconds) of top-$p$ sampling of Llama3 operator for a single batch. The baseline is labeled PyTorch and supports only discrete distributions with at most $2^{24}$ elements.
[…]
[Execution time (in milliseconds) of top-$p$ sampling of Llama3 operator for a single batch.]Execution time (in milliseconds) of top-$p$ sampling of Llama3 operator for a single batch.
Figure 13 depicts the execution time of drawing one sample using the top-$p$ sampler as it is implemented in the Llama3 model, see (Meta AI, 2024). PyTorch corresponds to the Ascend implementation using the baseline sort and cumsum operators. The lineplots with labels $s=32,64,128$ are similar to the baseline by replacing the sort() and cumsum() operators with the proposed radix sort and multi-core scan, respectively. The figure shows that the baseline top-$p$ sampling implementation scales poorly, mainly because the baseline torch.cumsum operator is not optimized for Ascend.
7. Conclusion & Future Work
We developed and evaluated efficient parallel scan algorithms tailored to the Ascend architecture by leveraging the power of the cube (matrix multiplication) unit. Our results demonstrate substantial performance improvements, with speedups ranging from 5×5$\times$5 × to 9.6×9.6$\times$9.6 × compared to vector-only implementations for sufficiently large input lengths. Additionally, we presented a multi-core Ascend scan algorithm that fully utilizes both the cube and vector units of Ascend, reaching up to 37.5% of the theoretical memory bandwidth.
Last, we extended our contributions to include crucial computational kernels for AI workloads, such as parallel split, compress/compact, and top-$p$ (nucleus) sampling, all exhibiting significant performance gains. Furthermore, our optimized implementation of radix sort, which utilizes matrix multiplications for parallel splits, showcases the potential of matrix engines in enhancing complex operations, offering up to $\times3$ speedup over the baseline.

### v2

@@ B.1 Multiple (Batched) Scans
Figure B.1: Batched scan algorithm based on ScanU. Blue and green denote the input and output array, respectively.
@@
Let’s discuss, as an example, a particular case where ScanU is used as a building block. The batched algorithm uses the same principles as the Algorithm 1 but also considers the 2-to-1 ratio between the vector and cube cores in the split Ascend architecture (910B). Figure B.1 depicts the main ideas behind our batched scan algorithm in this case. The algorithm starts by computing the local scans of size $s$ of $\bm{x}$ of all input arrays. Each cube core computes the local scans of a tile of size $\ell$ in two batches at the same time; once the tiles of the first two batches are ready, two distinct vector cores will complete the scans independently over each batch by propagating the partial sums within the tiles. This process is pipelined through AscendC to efficiently use all available hardware (vector) resources. The second batched scan algorithm extends ScanUL1 (Algorithm 2) so that each AI core computes a scan on a separate array in the batch.
@@
Figure B.2: Execution time ratio between ScanUL1 and ScanU batched scan algorithms for various array length ($x$-axis) and batch sizes ($y$-axis). Baseline is ScanU.
@@
Figure B.2 compares the two batched scan algorithms presented for increasing input batch size and array length. The first batched scan algorithm, based on ScanU (Algorithm 1), is used as our reference/baseline. Both algorithms have the same tiling strategy based on the input shapes to ensure a fair comparison. The figure demonstrates that the algorithms perform well in different cases and, more importantly, complement each other. In particular, ScanU is superior when the batch size is greater than $18$, and the input length is smaller than $4K$. On the other hand, ScanUL1 is superior when the batch size is smaller than $18$ and the input length larger than $4K$.
B.2 Exclusive and int8 scan
@@
Here, we discuss a few extensions of the multi-core scan algorithm that we have implemented. In particular, we added support for exclusive scans and integer inputs. Typically, AI accelerators support low-precision arithmetic since inference of deep learning models is robust to extreme levels of quantization [36]. In particular, Ascend supports input matrices of 8-bit integers with output/accumulation in 32-bit integers. Since scan is a memory-bound operator, there is an opportunity to improve performance in terms of elements per second processed; see Figure B.3. We have implemented a specialization of Algorithm 3 for integers with 8 bits, which is extensively used in our scan applications.
@@
We implemented exclusive scan by writing the output of the inclusive scan to global memory shifted by one element, discarding the last value and writing zero to the first position using the first block.
B.3 Weighted Sampling
@@
We implement a parallel weighted sampling kernel using the well-known inverse transform sampling approach and a scan to compute the cumulative distribution. Given an array $\bm{w}$ of $n$ positive weights, the goal is to draw a sample with proportional probability to the weights. The output is an index $i$ of $\bm{w}$ with probability proportional to $w_{i}$. First, we scan $\bm{w}$, and then, given a uniform sample $\theta\in{[0,1]}$, we invoke the SplitInd kernel with input $\text{scan}(\bm{w})$ and the element-wise predicate $\theta*\sum_{i}{w_{i}}$. The last entry of the output indices array of SplitInd contains the weighted sample. For more advanced parallel weighted sampling techniques, see [19].
@@
The performance improvement of our proposed weighted sampling kernel is not significant compared to the baseline for a single sample. However, our implementation does provide a functional improvement compared to the baseline operator. Indeed, the baseline Ascend weighted sampling operator torch.multinomial supports discrete distributions with support size up to $2^{24}$ elements, whereas our approach can support distributions with arbitrary support size. We leave as future work any further possible improvements on parallel weighted sampling. In particular, for the multiple sample generation scenario, the parallel alias table construction of [19] seems to be a promising direction.
B.4 Scan on low bit-width data types
@@
Next, we investigate the additional performance benefits of taking advantage of the lower-precision input data (int8) capability of the cube unit. Figure B.3 depicts the performance of the multi-core scan algorithm in terms of giga elements per second for input data types float16 and $8$-bit integers (int8). As depicted, there is only a small performance improvement for relatively small length, which is mainly due to our current implementation. For example, additional performance can be achieved since the Cube unit can multiply, in a single instruction, matrices of sizes $128\times 256$ and $256\times 128$ in int8.
@@
Figure B.3: Giga elements per second comparison of MCScan for float16 (fp16) and 8-bit integers (int8) input data types.
B.5 Radix sort on low bit-width inputs
Here, we provide an estimate of the expected performance improvements of radix sort for the case where the input data type has $4$ or $8$ bit-width. To do so, we measure the execution time of our radix sort kernel by reducing the number of radix sort iterations from $16$ to $4$ or $8$ (algorithmic correctness is not guaranteed). Notice that we still read and write using 16 bits per element. Figure B.4 shows a significant performance improvement of roughly $1.5\times$ and $3\times$ of our radix sort kernel for $8$ and $4$ bits inputs, respectively.
Figure B.4: Radix sort opportunity analysis for low bit-width inputs.